Calculate the electric field strength 30 cm from a 5 nC charge.


1
Expert's answer
2020-07-14T08:51:56-0400

According to the Coulomb's law, the electric field strength of a point charge is given by:


"E = k\dfrac{q}{r^2}"

where "q = 5\times 10^9C" is the charge, "r = 30cm = 0.3m" is the distance to point at which we consider the field and "k = 9\times 10^9N\cdot m^2/C^2" is the constant.

Substituting the numerical values, obtain:

"E = 9\times 10^9\cdot \dfrac{5\times 10^{-9}}{0.3^2}=500V/m"

Answer. 500 V/m.